# Efficient simulation of the SABR model

### v1

@@ 3 First simulation step: Sampling average variance $I_{t}^{h}$
3.1 Sampling $I_{t}^{h}$ from the moment-matched SLN approximation
@@
where $X$ is a standard normal variate. Here, $\mu$ is the mean of $Y$, $\lambda$ and $\sigma^{2}$ are the weight and variance of the LN component, respectively. When $\lambda=1$, $Y$ is reduced to the LN variable with mean $\mu$ and log variance $\sigma^{2}$.
The mean, coefficient of variation, skewness, and ex-kurtosis of $Y$ are given by
$E(Y)=\mu,\quad\text{Cv}(Y)=\lambda\sqrt{w},\quad\text{Sk}(Y)=\sqrt{w}(w+3),\quad\text{and}\quad\text{Ek}(Y)=w(w^{3}+6w^{2}+15w+16),$
where $w=\exp\left(\sigma^{2}\right)-1\;(w>0)$.
@@
While the skewness and ex-kurtosis of distribution $Y$ follow the conventional definitions, the coefficient of variation is defined as the ratio of standard deviation to mean:
$\mathrm{Cv}(Y)=\frac{\sqrt{\mathrm{Var}(Y)}}{E(Y)}.$
The coefficient of variation is an important dimensionless quantity that characterizes non-negative distributions such as the LN or SLN distributions.
@@
Given mean $\mu>0$, coefficient of variance $v>0$, and skewness $s>0$, the parameters of the SLN distribution matching $\mu$, $v$, and $s$ are analytically given by
$\sigma=\sqrt{\ln(1+w)}\quad\text{for}\quad w=4\sinh^{2}\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)$
$\text{and}\quad\lambda=\frac{v}{2\sinh\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)}.$
@@
$s^{2}=w(w+3)^{2}\quad\text{for}\quad w=\exp\left(\sigma^{2}\right)-1.$
@@
The expression for $w$ follows from $\cosh\theta-1=2\sinh^{2}(\theta/2)$. Next, we fit $\lambda$ to the given coefficient of variance $v$:
$\lambda=\frac{v}{\sqrt{w}}=\frac{v}{2\sinh\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)}.\qed$
@@
Next, we derive the first four conditional moments of $I_{t}^{h}$, whose analytic formulas are presented in the next proposition.
@@
Given $\sigma_{t}$ and $\sigma_{t+h}$, the first four raw moments of $I_{t}^{h}$ (i.e., $\mu=E(I_{t}^{h})$ and $\mu^{\prime}_{k}=E\left((I_{t}^{h})^{k}\right)$) are obtained as below.
@@
$\mu^{\prime}_{2}$ $=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{2}\frac{1}{\hat{\nu}^{2}}\left[m_{2}-cm_{1}\right],$
@@
where the conditioning variable, $\sigma_{t+h}$ and $\hat{Z}$, are exchangeable by $\hat{Z}=\frac{1}{\hat{\nu}}\ln(\sigma_{t+h}/\sigma_{t})$, and the coefficients, $c$ and $m_{k}\,(k=1,2,\dots,4)$, are the functions of them:
$c(\hat{Z}):=\cosh(\hat{\nu}\hat{Z})=\frac{1}{2}\left(\frac{\sigma_{t+h}}{\sigma_{t}}+\frac{\sigma_{t}}{\sigma_{t+h}}\right)\quad\text{and}\quad m_{k}(\hat{Z})=\frac{N(\hat{Z}+k\hat{\nu})-N(\hat{Z}-k\hat{\nu})}{2k\hat{\nu}\;n\left(\sqrt{\hat{Z}^{2}+(k\hat{\nu})^{2}}\right)},$
@@
The analytic derivation is based on the conditional moments of the exponential functional of BM (Matsumoto and Yor, 2005, (5.4)). See A for the derivation. ∎
@@
$v=\frac{\sqrt{\mu_{2}}}{\mu},\quad s=\frac{\mu^{\prime}_{3}-3\mu\mu^{\prime}_{2}+2\mu^{3}}{\mu_{2}\sqrt{\mu_{2}}},\quad\text{and}\quad\kappa=\frac{\mu^{\prime}_{4}-4\mu\mu^{\prime}_{3}+6\mu^{2}\mu^{\prime}_{2}-3\mu^{4}}{\mu_{2}^{2}}-3,$
@@
From Propositions 1 and 2, it is possible to sample $I_{t}^{h}$ from the SLN variable whose parameters are fitted to the first three moments of $I_{t}^{h}$ exactly. However, we take a simpler computational approach that uses a fixed $\lambda$ obtained from the small-time limit. It is shown to be effective as well.
Proposition 3 (SLN in the small-time limit).
@@
Proof.
In A, the coefficient of variation and skewness of $I_{t}^{h}$ are expanded around $\hat{\nu}=0$ as
$v=\frac{\sqrt{\mu_{2}}}{\mu}=\frac{\hat{\nu}}{\sqrt{3}}+O(\hat{\nu}^{3})\quad\text{and}\quad s=\frac{\mu_{3}}{\mu_{2}\sqrt{\mu_{2}}}=\frac{6\sqrt{3}}{5}\hat{\nu}+O(\hat{\nu}^{3}).$
Using Taylor’s expansion, we have
so that the shift parameter $\lambda$ converges to
Once $\lambda$ is determined, we fit $\sigma$ to match the coefficient of variation:
As we demonstrate in the numerical results shown in Figure 1, this approximation is highly reliable not only for small $\hat{\nu}$ but also for $O(\hat{\nu})=1$. ∎
Figure 1: Variance (left), skewness (middle), and Ex-kurtosis (right) of $I_{t}^{h}$ as functions of $\hat{Z}$ for $\hat{\nu}=0.4$. We compare our SLN approximation with $\lambda=5/6$ (yellow), the SLN approximation fitted up to skewness (green), and the LN approximation fitted up to variance (blue) to the true values (dashed).
@@ 3.2 Comparison to other simulation methods in literature
Our sampling method for $I_{t}^{h}$ is much faster and easier to implement compared to other methods in the literature, such as Chen et al. (2012), Cai et al. (2017), Leitao et al. (2017a, b) and Cui et al. (2021).
@@
Our method is considered as an extension of Chen et al. (2012) as they use an LN approximation matched to mean and variance of $I_{t}^{h}$. However, they derive the mean and variance only in the small-time limit, so the time step should be small accordingly. We enhance the method by deriving the moments analytically and matching to higher moments with the SLN approximation. Figure 1 displays the accuracy of variance, skewness, and ex-kurtosis of $I_{t}^{h}$ obtained from our SLN approximation. The left figure shows that Chen et al. (2012)’s small-time approximation of the variance is not accurate across the dependency $Z$ while our approximation is exact. The middle figure shows that the SLN approximation gives the correct skewness value as the parameters are fitted up to skewness. Moreover, the skewness under the simple SLN scheme with $\lambda=5/6$ also agrees well for $\nu=0.2$. Even though the two SLN schemes do not use the ex-kurtosis in the parameter calibration, the ex-kurtosis is close to the true value, justifying our choice of the SLN random variable.
Cai et al. (2017) computed the analytic form of the Laplace transform of the CDF of the reciprocal of the average variance 1/$I_{t}^{h}$, then used the Euler inversion formula to obtain the CDF of 1/$I_{t}^{h}$. After obtaining the CDF, they used the inverse transform method, which requires some tedious root-finding procedures such as Newton’s method or bisection method to sample $I_{t}^{h}$.
@@
Leitao et al. (2017a, b) used recursion schemes to find the characteristic function of the logarithm of $I_{t}^{h}$ and then employed the inverse Fourier technique COS method (Fang and Oosterlee, 2008) to recover its PDF. They then computed the Pearson correlation coefficient of $\ln I_{t}^{h}$ and $\ln\sigma_{t+h}$, then the bivariate copula distribution based on the CDF of $\ln\sigma_{T}$ and approximate CDF of $\ln I_{t}^{h}$. Finally, they adopted the direct inverse transform method based on linear interpolation to sample $I_{t}^{h}$.
@@
Cui et al. (2021) used the continuous-time Markov chain (CTMC) to approximate the space of the volatility process with finite number of states. In their procedure, they first fixed the volatility grids, computed the transition probability density at these grids based on the exponential of the generator matrix and used the inverse transform method to sample the volatility. After obtaining the volatility values, they computed the conditional characteristic function of the integrated variance, again based on the exponential of the generator matrix. They then adopted the Fourier sampler to get the CDF of the integrated variance, which first computed the PDF of the integrated variance via the orthogonal projection. Finally, they simulated the integrated variance using the closed-form inverse function.
@@
Putting aside the issue of accuracy of the simulation methods of $I_{t}^{h}$, these simulation methods invariably include some tedious and time-consuming procedures such as the adoption of the inverse Laplace/Fourier transform calculation, root-finding steps in the inverse transform methods or computation of the exponential of matrices. However, our SLN simulation methods for $I_{t}^{h}$ exhibits easy and fast implementation. As shown in the numerical experiments reported in Section 5, our simulation methods for $I_{t}^{h}$ demonstrate significant savings in computational time and high level of accuracy.
@@ 4 Second simulation step: Sampling forward price $F_{t+h}$
This section presents our CEV approximation for $F_{t+h}$ and its sampling method.
4.1 CEV approximation of the conditional price
Proposition 4 proposes our approximation of $F_{t+h}$ for the general case. Our choice of the approximation is a very natural and intuitive extension of the special cases discussed in Section. 2.3. The only technical component is the choice of the conditional terminal forward price that we adopt from Remark 1.
Proposition 4 (CEV approximation of the conditional terminal forward price).
For $10<\beta<$ 1 and $-1\leq\rho\leq 1$, conditional on $\sigma_{t+h}$ and $I_{t}^{h}$ (and the filtration up to time $t$), we approximate $F_{t+h}$ as a CEV distribution
$F_{t+h}\,|\,\sigma_{t+h},I_{t}^{h}\;\sim\;\mathcal{CEV}_{\beta}(\bar{F}_{t}^{h},\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}),$ (10)

### v2

@@ 3 First simulation step: Sampling average variance $I_{t}^{h}$
3.1 Sampling $I_{t}^{h}$ from the moment-matched shifted lognormal (SLN) approximation
@@
where $X$ is a standard normal variate. Here, $\mu$ is the mean of $Y$, $\lambda$ and $\sigma^{2}$ are the shifted parameter and variance of the LN component, respectively. When $\lambda=1$, $Y$ is reduced to the LN variable with mean $\mu$ and log variance $\sigma^{2}$.
The coefficient of variation, skewness, and ex-kurtosis of $Y$ are given by
$v(Y)=\lambda\sqrt{w},\quad\text{s}(Y)=\sqrt{w}(w+3),\quad\text{and}\quad\text{k}(Y)=w(w^{3}+6w^{2}+15w+16),$
respectively, where $w=e^{\sigma^{2}}-1\;(w>0)$.
@@
While the skewness and ex-kurtosis of distribution $Y$ follow the conventional definitions, the coefficient of variation is defined as the ratio of the standard deviation to mean:
$v(Y)=\frac{\sqrt{\mathrm{Var}(Y)}}{E(Y)}.$
The coefficient of variation is an important dimensionless quantity that characterizes non-negative distributions such as the LN and SLN distributions.
@@
Given mean $\mu>0$, coefficient of variance $v>0$, and skewness $s>0$, the parameters of the SLN distribution matching $\mu$, $v$, and $s$ admit the following analytic formulas:
$\sigma=\sqrt{\ln(1+w)}\quad\text{for}\quad w=4\sinh^{2}\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right),\quad\text{and}\quad\lambda=\frac{v}{2\sinh\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)}.$
@@
$s^{2}=w(w+3)^{2}\quad\text{for}\quad w=e^{\sigma^{2}}-1.$
@@
The expression for $w$ follows from $\cosh\theta-1=2\sinh^{2}\frac{\theta}{2}$. Next, we fit $\lambda$ to the given coefficient of variance $v$:
$\lambda=\frac{v}{\sqrt{w}}=\frac{v}{2\sinh\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)}.$
@@
Next, we derive the first four conditional moments of $I_{t}^{h}$, whose analytic formulas are presented in Proposition 2.
@@
Given $\sigma_{t}$ and $\sigma_{t+h}$, the first four raw moments of $I_{t}^{h}$ (i.e., $\mu=E(I_{t}^{h})$ and $\mu^{\prime}_{k}=E\left((I_{t}^{h})^{k}\right)$) are given by
@@
$\mu^{\prime}_{2}$ $=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{2}\frac{1}{\hat{\nu}^{2}}\left(m_{2}-cm_{1}\right),$
@@
where the conditioning variable, $\sigma_{t+h}$ and $\hat{Z}$, are exchangeable by $\hat{Z}=\frac{1}{\hat{\nu}}\ln(\sigma_{t+h}/\sigma_{t})$, and the coefficients, $c$ and $m_{k}\,(k=1,2,3,4)$, are functions of $\sigma_{t}$, $\sigma_{t+h}$ and $\hat{Z}$:
$c(\hat{Z}):=\cosh\hat{\nu}\hat{Z}=\frac{1}{2}\left(\frac{\sigma_{t+h}}{\sigma_{t}}+\frac{\sigma_{t}}{\sigma_{t+h}}\right)\quad\text{and}\quad m_{k}(\hat{Z})=\frac{N(\hat{Z}+k\hat{\nu})-N(\hat{Z}-k\hat{\nu})}{2k\hat{\nu}\;n\left(\sqrt{\hat{Z}^{2}+(k\hat{\nu})^{2}}\right)},$
@@
The analytic derivation is based on the conditional moments of the exponential functional of Brownian motion (Matsumoto and Yor, 2005, (5.4)). See A for the derivation. ∎
@@
$v(I_{t}^{h})=\frac{\sqrt{\mu_{2}}}{\mu},\quad s(I_{t}^{h})=\frac{\mu^{\prime}_{3}-3\mu\mu^{\prime}_{2}+2\mu^{3}}{\mu_{2}\sqrt{\mu_{2}}},\quad\text{and}\quad\kappa(I_{t}^{h})=\frac{\mu^{\prime}_{4}-4\mu\mu^{\prime}_{3}+6\mu^{2}\mu^{\prime}_{2}-3\mu^{4}}{\mu_{2}^{2}}-3,$
@@
From Propositions 1 and 2, it is possible to sample $I_{t}^{h}$ from the SLN variable whose parameters are fitted to the first three moments of $I_{t}^{h}$ exactly. However, we take a much simpler computational approach that uses a fixed $\lambda$ obtained from the small-time limit. It is shown to be effective as well (see Algorithm 1). In A, the coefficient of variation and skewness of $I_{t}^{h}$ are expanded around $\hat{\nu}=0$ as
$v=\frac{\sqrt{\mu_{2}}}{\mu}=\frac{\hat{\nu}}{\sqrt{3}}+O(\hat{\nu}^{3})\quad\text{and}\quad s=\frac{\mu_{3}}{\mu_{2}\sqrt{\mu_{2}}}=\frac{6\sqrt{3}}{5}\hat{\nu}+O(\hat{\nu}^{3}),\quad\text{where }\hat{\nu}=\nu\sqrt{h}$
Using Taylor’s expansion, we obtain
so that the shift parameter $\lambda$ converges to
Once $\lambda$ is determined, we fit $\sigma$ to match the coefficient of variation $v$:
As we demonstrate in the numerical results shown in Figure 1, this approximation is highly reliable not only for small $\hat{\nu}$ but even for $O(\hat{\nu})=1$.
We summarize the above results into Algorithm 1 for approximating conditional $I_{t}^{h}$ as follows:
@@
Algorithm 1 Approximation of the conditional $I_{t}^{h}$
Figure 1: Comparison of accuracy of calculating the skewness (left) and ex-kurtosis (right) of $I_{t}^{h}$ as functions of $\hat{Z}$ for $\hat{\nu}=0.4$. We compare our SLN approximation with $\lambda=5/6$ (orange), the SLN approximation fitted up to skewness (green), and the LN approximation fitted up to variance (blue) to the true values (dashed).
@@ 3.2 Comparison to other simulation methods in literature
Our sampling method for the approximation of conditional $I_{t}^{h}$ is much faster and easier to implement compared to other methods in the literature, such as Chen et al. (2012), Cai et al. (2017), Leitao et al. (2017a, b) and Cui et al. (2021).
@@
Our method is considered as an extension of Chen et al. (2012). Their LN approximation is matched up to mean and variance of $I_{t}^{h}$, while our approximation is matched up to the first three moments. Furthermore, they derive the mean and variance only in the small-time limit, so the time step has to be chosen to be small accordingly. We enhance the method by deriving the moments analytically and matching to higher moments with the SLN approximation. Figure 1 displays accuracy of skewness and ex-kurtosis of $I_{t}^{h}$ obtained from our SLN approximation. The left figure shows that the SLN approximation (green curve) gives the correct skewness value across the dependency of $\hat{Z}$ since the parameters are fitted up to skewness. Moreover, the skewness under the simple SLN scheme with $\lambda=5/6$ (orange curve) also agrees well. Even though the two SLN schemes do not use ex-kurtosis in the parameter calibration, the SLN approximation of the ex-kurtosis of $I_{t}^{h}$ is close to the true value. This justifies that our chosen SLN approximation (either exact $\lambda$ or $\lambda=\frac{5}{6}$) is a good choice of approximation of $I_{t}^{h}$ (see right figure). As evidenced from the plots in Figure 1, the SLN approximation (either exact $\lambda$ or $\lambda=\frac{5}{6}$) outperforms the LN approximation.
[…]
@@
Leitao et al. (2017a, b) used recursion schemes to find the characteristic function of the logarithm of $I_{t}^{h}$ and then employed the inverse Fourier technique COS method (Fang and Oosterlee, 2008) to recover its PDF. Next they computed the Pearson correlation coefficient of $\ln I_{t}^{h}$ and $\ln\sigma_{t+h}$, then the bivariate copula distribution based on the CDF of $\ln\sigma_{T}$ and approximate CDF of $\ln I_{t}^{h}$. Finally, they adopted the direct inverse transform method based on linear interpolation to sample $I_{t}^{h}$.
@@
Cui et al. (2021) used the continuous-time Markov chain (CTMC) in the space of the volatility process with finite number of states. In their procedure, they first fixed the volatility grids, computed the transition probability density at these grids based on the exponential of the generator matrix and used the inverse transform method to sample the volatility. After obtaining the volatility values, they computed the conditional characteristic function of the integrated variance, again based on the exponential of the generator matrix. They then adopted the Fourier sampler to get the CDF of the integrated variance, by first computing the PDF of the integrated variance via the orthogonal projection. Finally, they simulated the integrated variance using the closed-form inverse function.
@@
Besides the issue of accuracy of the simulation of $I_{t}^{h}$, these simulation methods invariably include tedious and time-consuming procedures such as the adoption of the inverse Laplace/Fourier transform calculation, root-finding steps in the inverse transform methods or computation of the exponential of matrices. However, our SLN simulation methods for $I_{t}^{h}$ exhibits easy and fast implementation. As shown in the numerical experiments reported in Section 5, our simulation methods for approximating $I_{t}^{h}$ demonstrate significant savings in computational time and high level of accuracy.
@@ 4 Second simulation step: Sampling forward price $F_{t+h}$
This section presents our CEV approximation for the forward price $F_{t+h}$ and the sampling of the CEV distribution via the exact simulation scheme of the shifted Poisson mixture gamma variate.
4.1 CEV approximation of the conditional forward price
In the uncorrelated SABR model ($\rho=0$), we observe that the conditional forward price $F_{T}$ follows the CEV distribution. To construct our SABR simulation scheme, we propose the CEV approximation of the conditional $F_{T}$ under the general case of $\rho\neq 0$ and $0<\beta<1$. We employ the operator splitting technique to derive the CEV approximation. Recall that a similar operator splitting approach has been used in the design of the simulation schemes for the rough Heston model in Bayer and Breneis (2024).
We rewrite the SDEs for the SABR volatility model [See Eq. (1)] into the following form:
$\frac{\text{d}F_{t}}{F_{t}^{\beta}}=\sigma_{t}(\rho\,\text{d}Z_{t}+\sqrt{1-\rho^{2}}\,\text{d}Z_{t}^{\perp})\quad\text{and}\quad\frac{\text{d}\sigma_{t}}{\sigma_{t}}=\nu\,\text{d}Z_{t},$ (10)
where $Z_{t}$ and $Z_{t}^{\perp}$ are uncorrelated BMs. Similar to the operator splitting approximation method in Lileika and Mackevičius (2021), we split the above SDEs into two parts:
$\frac{\text{d}F_{t}^{(1)}}{[F_{t}^{(1)}]^{\beta}}=\rho\sigma_{t}\,\text{d}Z_{t}\quad\text{and}\quad\frac{\text{d}\sigma_{t}}{\sigma_{t}}=\nu\,\text{d}Z_{t}$ (11a)
and
$\frac{\text{d}F_{t}^{(2)}}{[F_{t}^{(2)}]^{\beta}}=\sqrt{1-\rho^{2}}\sigma_{t}\,\text{d}Z_{t}^{\perp}\quad\text{and}\quad\frac{\text{d}\sigma_{t}}{\sigma_{t}}=\nu\,\text{d}Z_{t}.$ (11b)
The operator splitting approximation method dictates that an approximation to the solution of $F_{t}$ in Eq. (10) with initial condition $F_{0}$ is given by $F_{t}\sim F_{t}^{(2)}\left(F_{t}^{(1)}(F_{0})\right)$, where $F_{t}^{(1)}(F_{0})$ is the solution of $F_{t}^{(1)}$ in Eq. (11a) with the initial condition $F_{0}^{(1)}=F_{0}$ and $F_{t}^{(2)}\left(F_{t}^{(1)}(F_{0})\right)$ is the solution of $F_{t}^{(2)}$ in Eq. (11b) with the initial condition $F_{0}^{(2)}=F_{t}^{(1)}(F_{0})$.
Next, we solve the simplified set of SDEs for $F_{t}^{(1)}$ and $F_{t}^{(2)}$. The SDEs in Eq. (11b) resembles the SABR model with $\rho=0$ so that $F_{t}^{(2)}|\,\sigma_{t},I_{0}^{t}$ follows the CEV distribution, where
$F_{t}^{(2)}|\,\sigma_{t},I_{0}^{t}\;\sim\;\mathcal{CEV}_{\beta}(F_{0}^{(2)},\rho_{\ast}^{2}\sigma_{0}^{2}hI_{0}^{t}),\quad\rho_{\ast}^{2}=1-\rho^{2}.$ (12)
However, there is no simple analytic solution to Eq. (11a) for $\rho\neq 0$ and $0<\beta<1$. We propose to employ the frozen coefficient approximation to approximate $F_{t}^{(1)}]^{\beta_{\ast}}$ by $F_{0}^{\beta_{\ast}}$ such that Eq. (11a) is approximated by
$\frac{\text{d}F_{t}^{(1)}}{F_{t}^{(1)}}=\frac{\rho\sigma_{t}}{[F_{t}^{(1)}]^{\beta_{\ast}}}\,\text{d}Z_{t}\approx\frac{\rho\sigma_{t}}{F_{0}^{\beta_{\ast}}}\,\text{d}Z_{t}.$ (13)
Let $F_{t}^{(a)}$ denote the solution of the above approximating SDE, which is considered as an approximation of $F_{t}^{(1)}$. By rewriting the associated SDE for $F_{t}^{(a)}$ into the form:
$\text{d}\ln F_{t}^{(a)}=\frac{\rho}{\nu F_{0}^{\beta_{\ast}}}\,\text{d}\sigma_{t}-\frac{\rho^{2}\sigma_{t}^{2}}{2F_{0}^{2\beta_{\ast}}}\,\text{d}t,$
the solution of $F_{t}^{(a)}$ can be easily found by direct integration, which gives
$F_{t}^{(a)}(F_{0})=F_{0}\exp\left(\frac{\rho}{\nu F_{0}^{\beta_{\ast}}}(\sigma_{t}-\sigma_{0})-\frac{\rho^{2}\sigma_{0}^{2}tI_{0}^{t}}{2F_{0}^{2\beta_{\ast}}}\right).$ (14)
By combining the above operator splitting procedure coupled with frozen coefficient approximation, the approximation solution of $F_{t}$ is obtained as $F_{t}^{(2)}\left(F_{t}^{(a)}(F_{0})\right)$. Conditional on $\sigma_{t}$ and $I_{0}^{t}$, we obtain the CEV approximation of the conditional terminal forward price $F_{t}$ in the SABR model as follows:
$F_{t}\,|\,\sigma_{t},I_{0}^{t}\;\sim\;\mathcal{CEV}_{\beta}(\bar{F}^{t}_{0},\rho_{\ast}^{2}\sigma_{0}^{2}tI_{0}^{t}),$ (15a)
where
$\bar{F}^{t}_{0}=F_{0}\exp\left(\frac{\rho}{\nu F_{0}^{\beta_{\ast}}}(\sigma_{t}-\sigma_{0})-\frac{\rho^{2}\sigma_{0}^{2}tI_{0}^{t}}{2F_{0}^{2\beta_{\ast}}}\right).$ (15b)
The above discussion considers the approximation of solution of $F_{t}$ with initial time zero and end time $t$. In our context, we solve for $F_{t+h}$ with initial time $t$ and end time $t+h$. Accordingly, we change the initial time zero to time $t$ (comparing Eqs. (15a, 15b) with Eqs. (16a, 16b)). The CEV approximation of the conditional forward price $F_{t+h}\,|\,\sigma_{t+h},I_{t}^{h}$ is summarized in Algorithm 2:
For $0<\beta<1$ and $1\leq\rho\leq 1$, conditional on $\sigma_{t+h}$ and $I_{t}^{h}$ (filtration up to time $t$), we approximate $F_{t+h}$ as a CEV distribution as follows:
$F_{t+h}\,|\,\sigma_{t+h},I_{t}^{h}\;\sim\;\mathcal{CEV}_{\beta}(\bar{F}_{t}^{h},\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}),$ (16a)